a motorist travels for one hour at an average speed of 72 km/h and then for two hours at an average speed of 90 km/h. Calculate his average speed for the whole journey
Found 2 solutions by Alan3354, sachi:
Answer by Alan3354(69443)

a motorist travels for one hour at an average speed of 72 km/h and then for two hours at an average speed of 90 km/h. Calculate his average speed for the whole journey
==============
Use d = r*t to get the total distance
Divide by the total time to get the average speed

Answer by sachi(548)

journey speed(km/hr) hrs distance(km)=speed*hrs
1st 72 1 72
2nd 90 2 180
total 3 252
so the average speed for the whole journey=252/3=84 km/hr
ans